OpenStudy (anonymous):

If 34.2 grams of lithium react with excess water, how many liters of hydrogen gas can be produced at 299 Kelvin and 1.21 atmospheres? Show all of the work used to solve this problem. 2 Li (s) + 2 H2O (l) yields 2 LiOH (aq) + H2 (g)

OpenStudy (abmon98):

Number of Moles=Mass(g)/Molar Mass(g/mol) 34.2/6.94=4.927 Moles of lithium Mole Ratio of Lithium to Hydrogen 2:1 2:1 4.927:x xmole of hydrogen=4.927*1/2=2.463 Use ideal gas equation PV=nrt Pressure is measured in pascals 1 atmosphere = 101,325 Pa 1.21 atmosphere=122603.5Pa--->1.21*101325 Volume is measured in m^3 The SI value for R is 8.31441 J K-1 mol-1. 122603.5*V=8.31441*2.463*299 V=0.0499 m^3 1m^3--->1000litre 0.0499m^-->49.9litre--->0.0499*1000/1

OpenStudy (anonymous):

im confused

OpenStudy (abmon98):

about which part?

OpenStudy (anonymous):

reading and understanding tha entire thing...

OpenStudy (abmon98):

moles Li = 34.2 g / 6.941 g/mol=4.927 the ratio between Li and H2 is 2 : 1 moles H2 = 4.927/2 = 2.463 to obtain the volume of hydrogen gas we use the equation PV=nrt P stands for pressure and is measured in pascal not in atmosphere thats why i changed the unit V is measured in m^3 but in your question Volume has to be litres so i changed units by * by 1000 n is 2.463 moles R is a gas constant 8.31441

OpenStudy (abmon98):

You use Pv=nrt if you have a gas and you are given P, V or T data, and asked to find moles, or P or V or T or molar mass of the gas.

OpenStudy (somy):

do u get it now? @ThaPaintballer good job @Abmon98 !!

OpenStudy (abmon98):

Thanks, @Somy :D

OpenStudy (anonymous):

kind of, still a little lost tho with tha conversions..

OpenStudy (somy):

well look when we use PV= nRT we need to use SI system units which are P= Pascal V= m3 n= mole R= 8.14 (this value is like this because of the units we are using --> http://www.csupomona.edu/~lllee/gasconstant.pdf ) T= Kelvin

OpenStudy (somy):

is it clear now? there is your way also meaning no conversions in that case it will be P= atm V= L n= mole R= 0.08206 (check out in the link i gave you-> you will see that unit will match the units im using here) T= Kelvin

OpenStudy (somy):

so you don't need to convert anything, u can just use it with your units BUT of course you gotta know when which R value to use depending on the units you are taking if you are taking P as atm and V as L then your R will be = 0.08206 if you are taking P as Pascal and V as m3 then your R = 8.31 (btw sorry i its 8.31 not 8.14 )

OpenStudy (somy):

does it make sense now?

OpenStudy (anonymous):

yeah i'm kind of starting to piece it together, so 8.31 is tha answer?

OpenStudy (somy):

no its R value - when u use V as m3 and P as Pa then u put the values into the formula and get the unknown V

OpenStudy (somy):

but since the question is asking for volume in Liters i'd suggest not to convert anything

OpenStudy (anonymous):

so in PV =nRT 8.31 is subsituted for V but how do i get tha other units?

OpenStudy (somy):

leave pressure as atm and use R as 0.08206

OpenStudy (somy):

nooo 8.31 is value of R!

OpenStudy (somy):

we do not know V and we need to find it

OpenStudy (somy):

we know P= 1.21atm and T= 299 Kelvin and R= 0.08206 and we also know mole of Hydrogen

OpenStudy (somy):

we do not know Volume so we can find it using PV= nRT make V as subject so ==> V= nRT/P

OpenStudy (anonymous):

okay, so what is 'n' replaced with?

OpenStudy (somy):

with mole

OpenStudy (anonymous):

umm ok... so its v = mole*0.08206*299 / 1.21?

OpenStudy (somy):

yes you have mole

OpenStudy (somy):

put mole also or you don't get how you got mole??

OpenStudy (anonymous):

yeah i didnt really understand when you said its mole. chem isnt exactly my subject haha

OpenStudy (anonymous):

oop, never mind i get it now. so tha final answer is 50 L? @Somy

OpenStudy (somy):

idk lol if you calculated the way i told you then yes :))

OpenStudy (somy):

yeah lol i just checked :D

OpenStudy (somy):

well done both of you @Abmon98 @ThaPaintballer :D

OpenStudy (anonymous):

thanks haha (: so which one of you two do i give tha medal to? :3

OpenStudy (somy):

him :) he did the whole job lol i didn't do much :)

OpenStudy (anonymous):

alright, thanks again both of you! (:

OpenStudy (somy):

:))

OpenStudy (abmon98):

@Somy your good at teaching :) i think somy should receive a medal because she made it clear. thank you btw and its my pleasure @thapaintballer